OpenStudy (anonymous):

For f(x)=2x+1 and g(x)=x^2-7 find (g/f) (x) answers A. x^2 -7/2x+1 B. 2x+1/x^2-7 = + squared 7 (= has a line through it, and + has line under it) C. x^2 -7/ 2x+1, x = -1/2 (= sign has line through it) D. 2x+1/x^2-7

OpenStudy (triciaal):

as before (g/f)(x) = g(x) divided by f(x) g(x) = x^2 -7 and f(x) = 2x + 1 g(x) / f(x) = (x^2 -7) / (2x + 1)

OpenStudy (triciaal):

any questions why it is option C?

OpenStudy (anonymous):

No I think I got how to find it.

OpenStudy (michele_laino):

please note that, by definition, is: \[(g/f)(x)=\frac{ g(x) }{ f(x) }=...\]

OpenStudy (anonymous):

ok thanks!! @Michele_Laino

OpenStudy (michele_laino):

thanks! @Tallan